Search for Z′ in the dimuon spectrum of B → K μμ

Sascha Liechti

University of Zurich - LHCb group

Overview

1. Decay

2. Search 3. Z′ cusp 4. Outlook

Decay

• Anomalies gave rise to new search for possible lepton

flavor universality violation

• Different coupling to leptons of different generations

▷ Most to τ, least to e

▷ How can we observe different generations at once?

Search

• According to SM the τ contribution should not be visible in

the dimuon spectrum

• Dependent on Wilson coefficients (C7), C9, C10

▷ Some SUSY models predict an amplification of the τ

contribution:

▷ Amplification due to new channel ▷ Amplification factor: 10 - 1000

⇒ τ contribution should become visible

• Cusp like shape in between J/Ψ and Ψ(2S) resonances

• At the moment still a bifurkated gaussian (shape will slightly change as we get the exact shape)

Search

3

Search

Extremely sensitive to cusp amplitude

Cusp amplitude: 6 · 10^{−7}  Cusp amplitude: 4 · 10^{−7}

What can we do for now:

• Generate toys (equivalent size of data taken in run I and II)

• Running model of the dimuon spectrum and the relevant

contributions for this search

• Rare nonresonant

• J/Ψ

• Ψ(2S)

• Cusp

• Will be added in the future: cc, B → DDK , DD^{∗}K , D^{∗}D^{∗}K

• Binned fit pdf to the data and calculate likelihood Next steps:

• Integrate true shape of the cusp

• Add cc, B → DDK , DD^{∗}K , D^{∗}D^{∗}K ▷ Only B → DD^{∗}K nonresonant

• Implement additional crosschecks

• Low energy regime fit:

▷ Improve sensitivity on charm part and tau tail by including a constraint from low q^{2}

7